# Understanding Long Videos in One Multimodal Language Model Pass

### v1

@@ 4 Multimodal Video Understanding Framework
We propose a Multimodal Video Understanding Framework that integrates multimodal information from video using natural language. Our approach adapts a pre-trained VLM to construct a powerful long video understanding agent without performing any additional training. We first introduce a VLM-based Frame Selection module, based on a pre-trained VLM and Likelihood Selection, to extract 8 frames per input video, optionally considering the task questions and candidate answers. We then extract three object-centric modalities from the selected frames and use a fixed template to aggregate them into natural language. Finally, by applying likelihood selection again, we successfully turn a pre-trained generative VLM into a discriminator for video-language-based selection tasks. An overview of our framework is illustrated in Fig. 3.
4.1 Frame Selection
Due to the redundancy between video frames, long videos are generally sub-sampled to lower frame rates before processing to solve video tasks. Selecting frames (from a single video) most suitable for each task has been explored under numerous settings [7], with recent works also utilizing LLMs [78]. Motivated by [78], we extend our likelihood selection strategy to this frame selection task, demonstrating the adaptability of our selection strategy for alternate use cases. Specifically, we start from the conditional likelihood of a VLM, noted as $P(y|x_{t},x_{v}^{l})$ where $x_{v}^{l}$ stands for the $l^{\mathrm{th}}$ frame sampled from a video $x_{v}$. We first uniformly downsample the raw video in temporal to $L=16$ frames. For a fixed question $x_{t}$ and target $y$ (details in Appendix 0.B), we iterate through all 16 frames $x_{v}^{l}$ extracted from $x_{v}$ to compare the conditional likelihoods. Finally, we sample the first $8$ images per video in descending order by the conditional likelihood.
@@ 4.1 Frame Selection
4.2 Object-Centric Modalities
Having established a preliminary framework for frame selection and LLM-based answer selection, we replace the LLM with a VLM to process an additional single most relevant frame, $x_{v}^{\tilde{l}}$ ($\tilde{l}$ is the index of the top match frame from frame selection). We then explore how actual video modality information can be leveraged for further improvement. To this end, we focus on object-centric modalities, considering all object categories across the video, spatial locations of individual object instances, and their movement across time. These three forms of information can be extracted from a given video using off-the-shelf pre-trained models.
@@ 4.2 Object-Centric Modalities
    Global Object Information In this stage, we introduce global information that spans beyond a single video frame. For each of the 8 selected frames, we utilize a generative VLM to create object lists and obtain a set of distinct object categories contained across the video.
@@
    Object Spatial Location Given objects present per frame, we utilize an open-vocabulary object detector to localize each object onto frame coordinates. Additionally, we utilize the similarity of feature vectors (obtained from the object detector) for the same class objects to track object instances across frames. Following prior work [52], we calculate the average center coordinates and scale value for each object instance across all frames. This results in a set of distinct objects $O$ across the video, $O=\{(o_{1},q_{1}),(o_{2},q_{2}),...\}$.  Here, $o_{k}$ describes the object category in natural language while $q_{k}$ contains the x, y coordinates of the object centre and the scale term (area of minimal object bounding box as a ratio to image size, i.e. box area $\div$ image size).
@@
    Object Motion Trajectory: Next, we leverage the calculated cross-frame object tracks and compute motion trajectories for each object. This modifies our set of distinct objects, pairing each object $o_{k}$ with its trajectory ($o_{k}^{1}\rightarrow o_{k}^{2}\rightarrow...$) across the video frames. We construct an updated set $Z=\{(o_{1},q_{1}^{1}\rightarrow q_{1}^{2}\rightarrow...),(o_{2},q_{2}^{1}\rightarrow q_{2}^{2}\rightarrow...),...\}$. Intuitively, this information should explicitly capture object motion information.
@@ 4.3 Language based Fusion
Inspired by [81], we construct our overall framework by injecting these three forms of object-centric information into our setup using natural language. We represent each modality in a fixed template-based fusion. Global object information is represented as a list of category labels, e.g., $x_{\mathrm{GOI}}=$person, oven, dishwasher, …, sink}. Object spatial location modifies this list to include center coordinates ($x,y$) and scale ($s$) where scale is the area percentage occupied by the best-fitting object bounding box. For e.g., $x_{\text{OSL}}=$person located at (0.2, 0.3, 0.07), … , oven located at (0.8, 0.6, 0.04)}. Finally, object motion trajectories update the list to contain frame-level trajectories, e.g., $x_{\text{OMT}}=$person moving as [0.2, 0.3, 0.07] $\rightarrow$ [0.2, 0.4, 0.06] $\rightarrow$ [0.2, 0.6, 0.08], oven moving as … }. Similar to the examples, information from each object-centric modality is represented in textual form to allow their direct fusion and integration into our framework (as additional language inputs). Therein, we describe the resulting setup, our overall framework MVU as follows,
$\hat{y}=\mathcal{F}_{\text{MVU}}(x_{t},x_{v}^{\tilde{l}},x_{\text{GOI}},x_{\text{OSL}},x_{\text{OMT}})$ (5)
We refer the reader to Appendix 0.A for further details on prompting and template operations used as well as examples for our language-based fusion.
4.4 Modality Constrained Variants
We next introduce two modality-constrained variants of our framework tagged Just LLM and Single Frame VLM (see Fig. 1). The former utilizes only the task question injected as a language ($x_{t}$) with no other task-specific information. Note how this variant does not access any information extracted from video. The latter utilizes an additional center visual frame ($x_{v}^{c}$), extracted from the center of the video ($x_{v}$) timeline. This variant accesses no video-specific data (e.g. temporal or motion information). The center frame usage ensures no temporal information leakage in frame selection.
We hypothesize that Just LLM with no access to task-specific knowledge is constrained to generate predictions utilizing its internal world knowledge (e.g. physics, culture, human common sense). We refer to this as world modality. In the case of Single Frame VLM, it has task information but is limited to a single frame, which could possibly provide important scene context. Therein, we refer to this variant as operating in the world and contextual modality information.
Further exploration of these framework variants is presented in Sec. 5.3.
@@ 5 Experiments
Models: Our framework utilizes three distinct off-the-shelf models for its various operations, namely a) an LLM / VLM for likelihood selection (including visual frame selection and textual answer selection), b) a generative VLM for extracting object list from a frame, and c) an open-vocabulary detector for object localization. We use LLaVA-v1.5-13B [32] for likelihood selection, frame selection, and frame object list generation. For object localization, we use OWL-ViT-B/32 [38]. Unless explicitly specified, we use the above setup in all our experiments. Variants of our framework uses LLMs Llama-2-7b-Chat, Gemma-7b-IT, and Mistral-7B-Instruct (default) for likelihood selection. Apart from these off-the-shelf models, our framework involves zero additional training. Also, note that no components of our framework undergo any form of video-level training.
Datasets: We use multiple video datasets for evaluation under question-answering or n-way classification settings. For video question answering, we select two datasets focused on long-form videos: EgoSchema [35], NExT-QA [65]. EgoSchema is a long-form ego-centric video question-answering benchmark, consisting of a 500-video public subset (EgoSchema-S) and a full 5000+ video evaluation set (EgoSchema-F) accessed only through evaluation servers. This dataset spans over 250 hours and is specially constructed to ensure that questions require awareness of a longer temporal window for correctly answering [35]. Example images of EgoSchema are shown in Fig. 4. NExT-QA similarly contains long-form videos with a focus on requiring causal & temporal action reasoning as well as common scene comprehension for correctly answering. It contains a validation set (NextQA-V) of 4996 video-questions pairs and a test set (NextQA-T) of 8564 video-question pairs. We also use a series of robotics datasets from the Open X-Embodiment robotics dataset [40] for video question answering in a different domain (more detail in Sec. 5.2). In our ablations, we utilize a fine-grained action recognition dataset, Something-Something-v2 [15], that contains 174 action categories focusing on object motions by replacing object category nouns with ‘something’ in textual descriptions of each action category.
Compute: Our framework requires no re-training of off-the-shelf models. Our evaluations utilize two NVIDIA RTX A5000 24GB GPUs for inference.
@@ 5.1 Long Video Question Answering
Table 1: Ego-Schema Dataset Evaluation: We report top-1 accuracy (%) for video question answering on the open-subset (500 videos) and closed full-set (8564 videos) of Ego-Schema [35]. Our proposed MVU achieves state-of-the-art performance on this benchmark under zero-shot operation with no video level training. We also draw attention to our modality-constrained baselines, Vid-LLM and Vid-VLM, that achieve surprisingly competitive performance on this benchmark. Approaches utilizing proprietary LLMs with significantly higher parameters (e.g. GPT-4) are de-emphasized.
@@
ImageViT[43] 1B 16 40.8 30.9
ShortViViT[43] 1B 16 47.9
SeViLA+ShortViViT[43] 5B 32 49.6 31.3
LongViViT[43] 1B 256 56.8 33.3
MC-ViT-L [3] 424M 128 62.6 44.4
FrozenBiLM[73] 1.2B 90 - 26.9
VIOLET[14] 198M 5 - 19.9
mPLUG-Owl[75] 7.2B 5 - 31.1
InternVideo[62] 478M 90 32.1
SeViLA[78] 4B 32 25.7 22.7
LLoVi[83] 13B 180 50.8 33.5
Vamos [61] 13B - - 36.7
LLoVi[83] 1.8T 180 58.3 50.3
Vamos [61] 1.8T - - 48.3
Just LLM (ours) 7B 0 45.8 29.1
Single Frame VLM (ours) 13B 1 55.8 36.4
@@
Long video question answering benchmarks aim to measure the causal and temporal reasoning [65] abilities of models over long temporal windows [35]. In this section, we evaluate our framework on two benchmark datasets and present our results in Table 1 and Table 2.
On EgoSchema dataset, results reported in Tab. 1 demonstrate the state-of-the-art performance of our framework in comparison to work operating under similar settings. We reiterate how our approach is both zero-shot and requires no video-level training. In contrast, multiple prior zero-shot works utilize pseudo-labels on the downstream task videos for transductive training [14, 75, 78]. We also highlight how our MVU even outperforms multiple recent supervised (video-caption) approaches [43]. Among zero-shot methods, while our results are subpar to the trillion parameter variants of [83, 61] that utilize proprietary closed language models, we outperform with a clear margin against the fairer comparison to their similar parameter variants (more in Appendix 0.D). We also draw attention to our two modality-constrained model variants, Just LLM and Single Frame VLM that achieve surprisingly competitive results on this task, even outperforming multiple recent works. Motivated by these results, we conduct further experimentation under modality constrained setting in Sec. 5.3.
Table 2: Next-QA Dataset Evaluation: We report top-1 accuracy (%) for the test-split of Next-QA dataset [65]. Our proposed MVU achieves state-of-the-art results under zero-shot settings with no video-level training. We also highlight our competitive performance across the dataset’s causal, why, and how sub-categories. In table header, ZS stands for zero-shot and VT stands for video level training.
@@
CoVGT [68] 149M 58.8 57.4 69.3 60.0
SeViT [23] 215M - - - 60.6
HiTeA [74] 297M 62.4 58.3 75.6 63.1
InternVideo [62] 478M 62.5 58.5 75.8 63.2
MC-ViT-L [3] 424M - - - 65.0
BLIP-2 [29] 4B 70.1 65.2 80.1 70.1
SeViLA [78] 4B 74.2 69.4 81.3 73.8
LLama-VQA-7B [24] 7B 72.7 69.2 75.8 72.0
Vamos [61] 7B 72.6 69.6 78.0 72.5
Just-Ask [72] 66M 31.8 30.4 36.0 38.4
VFC [39] 164M 45.4 51.6 64.1 51.5
InternVideo [62] 478M 43.4 48.0 65.1 49.1
SeViLA[78] 4B 61.3 61.5 75.6 63.6
CaKE-LM w HGA [57] 2.7B 35.3 34.7 36.9 34.8
CaKE-LM w CoMem [57] 2.7B 35.7 35.3 36.8 34.9
LLoVi [83] 13B 55.6 47.9 63.2 54.3
MVU (ours) 13B 55.4 48.1 64.1 55.2
We next evaluate our framework on the NextQA benchmark and report these results in Tab. 2. Our MVU achieves competitive results under zero-shot settings. While [78] outperforms our approach, we note how they require video-caption localization pretraining and appears to overfit to this dataset considering their relatively lower performance on other datasets (see Tab. 1). Considering the zero-shot, video training-free aspect of our proposed MVU, we achieve state-of-the-art results in comparison to similar prior work.
Table 3: Robotics Domain Action Recognition Evaluation: We report top-1 accuracy (%) for our VideoQA style action recognition formulation (details in Sec. 5.2) on the Open X-Embodiment collection of datasets [40]. Our proposed MVU achieves clear improvements over random selection and a single-frame baseline. In the table header, Obs. (observation), size, CC (class count) stand for the camera used, number of videos, and number of unique language instructions per dataset, respectively. In the observation column, T stands for third-person view, a stationary camera that does not move with the robot. While F means first-person view where the camera is mounted on the moving robot. SF-VLM stands for Single Frame VLM.
Dataset Obs. Size CC Random SF-VLM MVU
ASU TableTop Manipulation [87, 86] T 110 83 13.6 20.9 20.9
Berkeley MVP Data [47] F 480 6 20.0 22.9 33.1
Berkeley RPT Data [46] F 908 4 24.6 20.6 20.6
CMU Play Fusion [8] T 576 44 20.3 34.2 35.6
CMU Stretch [2, 37] T 135 5 23.0 18.5 24.4
Furniture Bench [19] T 5100 9 20.2 12.1 16.1
Furniture Bench [19] F 5100 9 20.2 13.1 12.7
CMU Franka Pick-Insert Data [55] T 631 7 18.7 18.9 21.2
CMU Franka Pick-Insert Data [55] F 631 7 23.1 59.6 49.3
Imperial F Cam T 170 17 20.0 24.7 24.1
Imperial F Cam F 170 17 23.5 23.5 20.6
USC Jaco Play [12] T 1085 89 21.8 16.2 18.0
USC Jaco Play [12] F 1085 89 19.4 28.0 32.4
NYU ROT [16] T 14 12 21.4 50.0 57.1
Roboturk [34] T 1959 3 34.7 43.1 37.9
Stanford HYDRA [4] T 570 3 35.1 44.6 68.2
Stanford HYDRA [4] F 570 3 31.2 44.0 48.9
Freiburg Franka Play [53, 36] F 3603 406 20.4 32.2 31.6
Freiburg Franka Play [53, 36] T 3603 406 19.7 21.8 24.0
LSMO Dataset [42] T 50 2 34.0 50.0 72.0
UCSD Kitchen [71] T 150 8 19.3 32.0 32.7
Austin VIOLA [88] T 150 3 26.7 33.3 33.3
Austin VIOLA [88] F 150 3 30.0 33.3 33.3
Total - 27000 - 22.1 23.5 25.0

### v2

@@ 4 Multimodal Video Understanding Framework
In this section, we introduce in detail our Multimodal Video Understanding (MVU) framework that integrates several information modalities extracted from video using natural language as a medium for information fusion. Our approach adapts off-the-shelf vision tools to construct a powerful long video understanding agent that requires no additional training on videos. We first utilize vision tools to extract information relevant to three object-centric modalities from uniformly sampled video frames. Next we leverage suitable prompt templates to aggregate these as natural language. This video level information is injected to our Single-Frame-VLM variant providing it with video specific awareness. We illustrate an overview of our framework in Figure 3.
4.1 Vision Tools for Video Analysis
Image trained VLMs contain information valuable for video tasks and have been widely used in prior work (Zhang et al., 2023a). In our proposed framework, we take a step further, exploring more off-the-shelf vision tools trained only on images, in particular object detection and object tracking approaches, in addition to a VLM re-purposed as an image captioner.
We use an image captioner to identify all unique objects present within a video. For this purpose, we prompt a generative vision language model to list all objects within a given video frame (image) in an open-ended manner. We note how a VLM trained only on images is sufficient for this. In our case, we use a VLM identical to the one in Zhang et al. (2023a) but applied on significantly less video frames, making our comparisons fair in terms of model size.
For the case of object detection, we use an open-vocabulary object detector from Minderer et al. (2022) that is trained only on images, and apply it with object category names from captioner to obtain their location information, i.e. image-space coordinates for each unique object. Given the lightweight nature of this detector in comparison to the image captioner, we note how it can be applied more densely (i.e. on more frames) than the captioner without increasing compute demand significantly. Furthermore, the detector acts as a secondary check, grounding the object category names to individual frames, and therein countering any object hallucinations by the captioner.
Our final tool is an object tracker from Wang et al. (2018) used to convert our per-frame object detections into motion trajectories spread across the entire video. We feed the tracking algorithm with the locations of each object alongside per-object features extracted from our detector in order to construct motion trajectories for each unique object.
@@ 4.1 Vision Tools for Video Analysis
4.2 Object-Centric Information Modalities
Given off-the-shelf tools suitable for extracting information from videos, we next focus on the exact forms of information, i.e. three object-centric information modalities. We consider all object categories across the video, spatial locations of individual object instances, and their movement across time. We define these as follows:
@@ 4.2 Object-Centric Information Modalities
    Global Object Information ($x_{\mathrm{GOI}}$): In this stage, we introduce global information that spans beyond a single video frame. For a given video, we first uniformly sample 8 frames. For each of the 8 selected frames, we utilize our image captioner to generate object lists and obtain a set of distinct object categories contained within each frame across the video.
@@
    Object Spatial Location ($x_{\mathrm{OSL}}$): Given objects present per video, we utilize our open-vocabulary object detector to localize each object category (from previous stage) onto frame coordinates. Categories not localized by the detector are dropped. Additionally, we utilize similarity of feature vectors for same class objects to track object instances across frames using our tracker. Following prior work (Ranasinghe et al., 2024), we calculate average center coordinates and scale value for each object instance across all frames. This results in a set of distinct objects $O$ across the video, $O=\{(o_{1},q_{1}),(o_{2},q_{2}),...\}$.  Here, $o_{k}$ describes the object category in natural language while $q_{k}$ contains the x, y coordinates of object centre and the scale term (area of minimal object bounding box as a ratio to image size, i.e. box area $\div$ image size).
@@
    Object Motion Trajectory ($x_{\mathrm{OMT}}$): Next, we leverage the calculated cross-frame object tracks and compute motion trajectories for each object. This modifies our set of distinct objects, pairing each object $o_{k}$ with its trajectory ($o_{k}^{1}\rightarrow o_{k}^{2}\rightarrow...$) across the video frames. We construct an updated set $Z=\{(o_{1},q_{1}^{1}\rightarrow q_{1}^{2}\rightarrow...),(o_{2},q_{2}^{1}\rightarrow q_{2}^{2}\rightarrow...),...\}$. Intuitively, this information should explicitly capture object motion information.
We provide further details including examples of each information modality for selected samples (video question pairs) in Appendix A.
This pipeline for extracting per-frame information using an image-trained VLM closely resembles prior work such as (Zhang et al., 2023a). While motivated by such work, we explore the direction of how more fine-grained information could be extracted from videos to solve these tasks more efficiently. Given the role of object interactions in defining the various actions and events in videos, we hypothesize that extracting object-centric information followed by explicit capturing of their temporal dependencies (as opposed to generic frame-level descriptions) would be highly beneficial to solving video tasks.
We also highlight how the use of vision tools to extract specific forms of information provides an additional level of interpretability to the functions of our overall pipeline. We next discuss how we fuse each of these information modalities using natural language as a medium.
@@ 4.3 Language based Fusion
Inspired by Zeng et al. (2022), we construct our overall framework by injecting these three forms of object-centric information into our setup using natural language. We represent each modality in a fixed template-based fusion. Global object information is represented as a list of category labels, e.g., $x_{\mathrm{GOI}}=$person, oven, dishwasher, …, sink}. Object spatial location modifies this list to include center coordinates ($x,y$) and scale ($s$) where scale is the area percentage occupied by the best-fitting object bounding box. For e.g., $x_{\text{OSL}}=$person located at (0.2, 0.3, 0.07), … , oven located at (0.8, 0.6, 0.04)}. Finally, object motion trajectories update the list to contain frame-level trajectories, e.g., $x_{\text{OMT}}=$person moving as [0.2, 0.3, 0.07] $\rightarrow$ [0.2, 0.4, 0.06] $\rightarrow$ [0.2, 0.6, 0.08], oven moving as … }. Similar to the examples, information from each object-centric modality is represented in textual form to allow their direct fusion and integration into our framework (as additional language inputs). Therein, we describe the resulting setup, our overall framework MVU as follows,
$\hat{y}=\mathcal{F}_{\text{MVU}}(x_{t},{x}_{v}^{c},x_{\text{GOI}},x_{\text{OSL}},x_{\text{OMT}})$ (1)
where ${x}_{v}^{c}$ is the center frame extracted from the video $x_{v}$. We refer the reader to Appendix A for further details on prompting and template operations used as well as examples for our language-based fusion.
@@ 5 Experiments
Experimental Setup: Our proposed MVU framework and its variants use off-the-shelf models trained on images, thus requiring no re-training of these models. For our evaluations, we directly use these models, utilizing two NVIDIA RTX A5000 24GB GPUs for inference. We evaluate on two video question answering datasets focused on long-form videos: EgoSchema (Mangalam et al., 2023) and NExT-QA (Xiao et al., 2021). We also evaluate using a series of robotics datasets from the Open X-Embodiment robotics dataset (Open-X-Embodiment-Collaboration et al., 2023) to test our model generality (more details in Section 5.2). We discuss further details of pretrained models and datasets in Appendix B. Also, note that none of the pretrained components of our framework undergo any form of video-level training.
@@ 5.1 Long Video Question Answering
Table 2: Ego-Schema Dataset Evaluation: We report top-1 accuracy (%) for video question answering on the open-subset (500 videos) and closed full-set (8564 videos) of Ego-Schema (Mangalam et al., 2023). Our proposed MVU achieves state-of-the-art performance on this benchmark under zero-shot operation with no video level training. We also draw attention to our modality-constrained baselines, Vid-LLM and Vid-VLM, that achieve surprisingly competitive performance on this benchmark. Approaches utilizing proprietary LLMs with significantly higher parameters (e.g. GPT-4) are de-emphasized.
@@
ImageViT (Papalampidi et al., 2023) 1B 16 40.8 30.9
ShortViViT (Papalampidi et al., 2023) 1B 16 47.9
SeViLA+ShortViViT (Papalampidi et al., 2023) 5B 32 49.6 31.3
LongViViT (Papalampidi et al., 2023) 1B 256 56.8 33.3
MC-ViT-L (Balavzevi’c et al., 2024) 424M 128 62.6 44.4
FrozenBiLM (Yang et al., 2022) 1.2B 90 - 26.9
VIOLET (Fu et al., 2022) 198M 5 - 19.9
mPLUG-Owl (Ye et al., 2023b) 7.2B 5 - 31.1
InternVideo (Wang et al., 2022) 478M 90 32.1
SeViLA (Yu et al., 2024) 4B 32 25.7 22.7
LLoVi (Zhang et al., 2023a) 13B 180 50.8 33.5
Vamos (Wang et al., 2023) 13B - - 36.7
LLoVi (Zhang et al., 2023a) 1.8T 180 58.3 50.3
Vamos (Wang et al., 2023) 1.8T - - 48.3
Just-LLM (ours) 7B 0 45.8 29.1
Single-Frame-VLM (ours) 13B 1 55.8 36.4
@@
Long video question answering benchmarks aim to measure causal and temporal reasoning abilities of models over long temporal windows (Xiao et al., 2021; Mangalam et al., 2023). In this section, we evaluate our framework on two benchmark datasets and present our results in Table 2 and Table 3.
On EgoSchema dataset, results reported in Table 2 demonstrate the state-of-the-art performance of our framework in comparison to work operating under similar settings. We reiterate how our approach is both zero-shot and requires no video-level training. In contrast, multiple prior zero-shot works utilize pseudo-labels on the downstream task videos for transductive training (Fu et al., 2022; Ye et al., 2023b; Yu et al., 2024). We also highlight how our MVU even outperforms multiple recent supervised (video-caption) approaches (Papalampidi et al., 2023). Among zero-shot methods, while our results are subpar to the trillion parameter variants of (Zhang et al., 2023a; Wang et al., 2023) that utilize proprietary closed language models, we outperform with a clear margin against the fairer comparison to their similar parameter variants (more in Appendix C).
Table 3: Next-QA Dataset Evaluation: We report top-1 accuracy (%) for the test-split of Next-QA dataset (Xiao et al., 2021). Our proposed MVU achieves state-of-the-art results under zero-shot settings with no video-level training. We also highlight our competitive performance across the dataset’s causal, why, and how sub-categories. In table header, ZS stands for zero-shot and VT stands for video level training.
@@
CoVGT (Xiao et al., 2023) 149M 58.8 57.4 69.3 60.0
SeViT (Kim et al., 2023) 215M - - - 60.6
HiTeA (Ye et al., 2023a) 297M 62.4 58.3 75.6 63.1
InternVideo (Wang et al., 2022) 478M 62.5 58.5 75.8 63.2
MC-ViT-L (Balavzevi’c et al., 2024) 424M - - - 65.0
BLIP-2 (Li et al., 2023a) 4B 70.1 65.2 80.1 70.1
SeViLA (Yu et al., 2024) 4B 74.2 69.4 81.3 73.8
LLama-VQA-7B (Ko et al., 2023) 7B 72.7 69.2 75.8 72.0
Vamos (Wang et al., 2023) 7B 72.6 69.6 78.0 72.5
Just-Ask (Yang et al., 2021) 66M 31.8 30.4 36.0 38.4
VFC (Momeni et al., 2023) 164M 45.4 51.6 64.1 51.5
InternVideo (Wang et al., 2022) 478M 43.4 48.0 65.1 49.1
SeViLA(Yu et al., 2024) 4B 61.3 61.5 75.6 63.6
CaKE-LM w HGA (Su et al., 2023) 2.7B 35.3 34.7 36.9 34.8
CaKE-LM w CoMem (Su et al., 2023) 2.7B 35.7 35.3 36.8 34.9
LLoVi (Zhang et al., 2023a) 13B 55.6 47.9 63.2 54.3
MVU (ours) 13B 55.7 48.2 64.2 55.4
We next evaluate our framework on the NextQA benchmark and report these results in Table 3. Our MVU achieves competitive results under zero-shot settings. While Yu et al. (2024) outperforms our approach, we note how they require video-caption localization pretraining and appears to overfit to this dataset considering their relatively lower performance on other datasets (see Table 2).
While these two datasets focus on MCQ style QnA, we also explore the generality of our MVU framework on open-ended style QnA tasks. We detail these evaluations in Appendix G, highlighting how MVU performance extends beyond MCQ style tasks.